# CLAMBER: A Benchmark of Identifying and Clarifying Ambiguous Information Needs in Large Language Models

### v1

@@ Abstract
[…]
@@ 1 Introduction
Given well-defined user queries, large language models (LLMs) have demonstrated remarkable proficiency in facilitating the information search process (Pan et al., 2023; Kamalloo et al., 2023; Zhang and Choi, 2023; Huang et al., 2024). They provide more precise search results with the help of the inherent knowledge stored within LLMs. Nonetheless, as evidenced by previous studies (Kuhn et al., 2023; Deng et al., 2023c, a), the practical utility of LLMs is hindered by unclear and ambiguous user queries in real-world scenarios. For instance, in a query like "what are the strategies for saving?", the term "saving" can have multiple interpretations, such as "saving money" or "saving from sins", depending on the user’s actual need. This necessitates LLMs proactively identifying (i.e., determine if the query is ambiguous or not) and clarifying the ambiguities rather than providing potentially incorrect answers that may not align with the user’s true needs, ultimately risking user trust and satisfaction (Deng et al., 2023b; Liao et al., 2023).
@@ 2 Related Works
Table 1: The proposed taxonomy of ambiguous queries and examples.
@@ 5.2 Fine-Grained Evaluation
Table 7: Overall ambiguity identification evaluation of LLMs with varying prompting schemes. ChatGPT emerges as the superior model to other open-sourced LLMs. We report BertScore (i.e., BS) and Help.
@@ Appendix E More Task Results
Table 8, 9, 10 present the results of all LLMs across different categories under three different settings: Zero-shot w/o CoT, Zero-shot w/ CoT, and Few-shot w/o CoT. We discover that while the exact values vary, the overall performance and analysis conclusions remain largely consistent with Sec 5.2.

### v2

@@ Abstract
$\star$\$$\star$\$footnotetext: Both authors contributed equally to this study.
@@ 1 Introduction
Given well-defined user queries, large language models (LLMs) have demonstrated remarkable proficiency in facilitating the information search process (Pan et al., 2023; Kamalloo et al., 2023; Zhang and Choi, 2023; Huang et al., 2023). They provide more precise search results with the help of the inherent knowledge stored within LLMs. Nonetheless, as evidenced by previous studies (Kuhn et al., 2023; Deng et al., 2023a), the practical utility of LLMs is hindered by unclear and ambiguous user queries in real-world scenarios. For instance, in a query like "what are the strategies for saving?", the term "saving" can have multiple interpretations, such as "saving money" or "saving from sins", depending on the user’s actual need. This necessitates LLMs proactively identifying (i.e., determine if the query is ambiguous or not) and clarifying the ambiguities rather than providing potentially incorrect answers that may not align with the user’s true needs, ultimately risking user trust and satisfaction (Liao et al., 2023).
@@
Table 1: The proposed taxonomy of ambiguous queries and examples. The clarifying questions of each example are provided in Table 8.
2 Related Works
@@ 5.2 Fine-Grained Evaluation
Table 7: Overall ambiguity clarification evaluation of LLMs with varying prompting schemes. ChatGPT emerges as the superior model to other open-sourced LLMs. We report BertScore (i.e., BS) and Help.
@@ C.7 Dolly Dataset
Category Query Clarifying Question
UNFAMILIAR Find the price of Samsung Chromecast. Do you mean Google Chromecast or Samsung Chromebook?
CONTRADICTION
The critic is in the restaurant.>X.
The butterfly is in the river.>Y.
The boar is in the theatre.>?
Is the category either human or indoor location?
LEXICAL Tell me about the source of Nile. Are you referring to the Nile river or the board game?
SEMANTIC When did he land on the moon? Who do you mean by "he" in the question?
WHO Suggest me some gifts for my mother. What are your mother’s interests, hobbies, or something she has been wanting?
WHEN How many goals did Argentina score in the World Cup? Which World Cup are you referring to?
WHERE Tell me how to reach New York. Please tell me where your departure place is?
WHAT Real name of gwen stacy in spiderman? Which one are you referring to: the actress,or the character?
Table 8: The example clarifying questions associated with ambiguous queries in Table 1. There are no discerning patterns according to the ambiguity category.
@@ Appendix E More Task Results
Table 9, 10, 11 present the results of all LLMs across different categories under three different settings: Zero-shot w/o CoT, Zero-shot w/ CoT, and Few-shot w/o CoT. We discover that while the exact values vary, the overall performance and analysis conclusions remain largely consistent with Sec 5.2.